OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3.

OpenStudy (anonymous):

g(3)=-13, so all u need to find out is f(-13)=-26+1=-25

OpenStudy (anonymous):

g(3) = -5*3 +2 = -13 and f(-13) = 2(-13)+1= -25

OpenStudy (joannablackwelder):

Do you understand why you start on the inside and work your way out?

OpenStudy (anonymous):

a little

OpenStudy (joannablackwelder):

You are wanting to find a composite function. So, g(x) is what you are plugging in for the x of f(x). You could find the function f(g(x)) and then plug in 3 into that function if that makes more sense to you.

OpenStudy (joannablackwelder):

f(g(x)) = 2(-5x+2)+1

OpenStudy (joannablackwelder):

Do you see how I did that?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

good work joanna!

OpenStudy (joannablackwelder):

When you substitute x=3 and simplify, you get -25. Same as the other folks, just a different method. Whatever works best for you. :)

OpenStudy (joannablackwelder):

Thanks @cody_123